# Topology Matters: Measuring Memory Leakage in Multi-Agent LLMs

### v1

@@ Abstract
Graph topology is a fundamental determinant of memory leakage in multi-agent LLM systems, yet its effects remain poorly quantified. We introduce MAMA (Multi-Agent Memory Attack), a framework that measures how network structure shapes leakage. MAMA operates on synthetic documents containing labeled Personally Identifiable Information (PII) entities, from which we generate sanitized task instructions. We execute a two-phase protocol: Engram (seeding private information into a target agent’s memory) and Resonance (multi-round interaction where an attacker attempts extraction). Over up to 10 interaction rounds, we quantify leakage as the fraction of ground-truth PII recovered from attacking agent outputs via exact matching. We systematically evaluate six common network topologies (fully connected, ring, chain, binary tree, star, and star-ring), varying agent counts $n\in\{4,5,6\}$, attacker-target placements, and base models. Our findings reveal consistent patterns: fully connected graphs exhibit maximum leakage while chains provide strongest protection; shorter attacker-target graph distance and higher target centrality significantly increase vulnerability; leakage rises sharply in early rounds before plateauing; model choice shifts absolute leakage rates but preserves topology rankings; temporal/locational PII attributes leak more readily than identity credentials or regulated identifiers. These results provide the first systematic mapping from architectural choices to measurable privacy risk, yielding actionable guidance: prefer sparse or hierarchical connectivity, maximize attacker-target separation, limit node degree and network radius, avoid shortcuts bypassing hubs, and implement topology-aware access controls.
@@
Jinbo Liuthanks: These authors are co-first authors.1   Defu Cao11footnotemark: 11   Yifei Weithanks: These authors contributed equally to data collection, analysis, and manuscript revision.1   Tianyao Su22footnotemark: 21 Yuan Liang22footnotemark: 21   Yushun Dong2   Yue Zhao1   Xiyang Hu3 1University of Southern California  2Florida State University  3Arizona State University {jinboliu,defucao,yifeiwei,tianyaos,yliang23,yue.z}@usc.edu yd24f@fsu.eduxiyanghu@asu.edu
@@ 1 Introduction
Multi-agent systems based on large language models (LLMs) are rapidly moving from prototype to real-world use (li2024llm_mas_survey). When viewed through the lens of network science, multi-agent LLM systems are distributed communication networks with message routing, shared memory, and coordination (kurose_ross; watts1998). This reveals a critical insight: these systems inherit structural vulnerabilities that govern information diffusion in communication networks, social systems, and epidemic models. Network topology—the pattern of connections between agents—becomes a first-order security parameter (newman_watts_science), where vulnerabilities at the topological layer can amplify local failures into system-wide compromises. Network science research demonstrate that structural properties—path length, node centrality, clustering coefficients, and long-range links—fundamentally govern how information spreads and how cascading failures occur (boccaletti2006). Recent work on LLM multi-agent security provides empirical support: connectivity patterns and inter-agent distances create pathways for adversarial diffusion (gsafeguard), with dense topologies like fully-connected graphs proving particularly vulnerable (netsafe; iclr_resilience).
Beyond the propagation of adversarial prompts or harmful content, the leakage of Personally Identifiable Information (PII) poses an equally critical threat in multi-agent architectures. Real-world deployments distribute confidential data—user credentials, API keys, proprietary tools, conversation histories—across agents through multi-turn interactions and role-specific memory. Recent investigations have found topology details, system prompts, and tool leaks in black-box configurations (masleak; safesearch), as well as co-hijacking caused by malicious input (codeexec; integrity). These findings underscore an urgent research gap: we lack systematic understanding of how communication topology shapes the leakage surface for private information in multi-turn, multi-agent interactions.
@@
Topology-aware security research has begun addressing these concerns through distance and connectivity analysis (netsafe), graph-level interventions (gsafeguard), and structural resilience comparisons (iclr_resilience). However, three critical gaps remain. First, prior work targets adversarial content propagation and task degradation (netsafe; gsafeguard) rather than fine-grained PII leakage dynamics. They lack metrics like "time to leak" or "success rate curves" for PII exposure. Second, data exfiltration studies do not systematically control for agent placement, graph distance, or interaction horizon (iclr_resilience; masleak), making it difficult to isolate topology effects. Third, while network theory predicts that sparse long-range links and shortcuts dramatically alter diffusion (watts1998; newman_watts_science), these structural phenomena have not been rigorously connected to information leakage in LLM-based systems.
@@
We introduce MAMA (Multi-Agent Memory Attack), a systematic framework for measuring how network topology governs memory leakage in multi-agent LLM systems. MAMA addresses the limitations of prior work through three key design principles. First, controlled synthesis: we generate synthetic task documents containing labeled PII entities spanning five categories (identity, contact, location, temporal, and regulated identifiers), then derive sanitized task instructions that contain no direct PII, ensuring any leakage originates from agent memory rather than task specification. Second, systematic topology variation: we evaluate six canonical graph structures (fully connected, ring, chain, binary tree, star, and star-ring) across team sizes ($n\in\{4,5,6\}$), systematically varying attacker-target placement to control for graph distance and node centrality (iclr_resilience). Third, temporal resolution: we execute a two-phase protocol—Engram (memory seeding) and Resonance (multi-round extraction)—tracking leakage dynamics over up to 10 interaction rounds. Our framework operationalizes network science concepts (distance, hubs, long-range links) as measurable security indicators (netsafe; gsafeguard), incorporating standardized attack prompts designed to elicit specific PII from target agents (codeexec) and fine-grained leakage metrics including attack success rate, time to first leak, shortest leakage path, and graded matching outcomes (exact, approximate, or trace-level) (safesearch).
@@
Our work makes four principal contributions. First, we develop a controllable experimental pipeline spanning data synthesis (documents with labeled PII), instruction sanitization (leak-free task generation), and a standardized two-phase interaction protocol, enabling systematic evaluation across graph families, team sizes, attacker-target configurations, and multi-round horizons. Second, we formalize a unified threat model with explicit agent roles (normal workers, target with private memory, internal attacker, optional reviewers/observers) and comprehensive leakage metrics (overall success rate, time to leak, shortest leakage paths, graded matching). Third, using MAMA we uncover consistent topology-leakage patterns: denser connectivity and shorter attacker-target distances increase leakage; fully connected and star-ring structures prove most vulnerable while chains and trees provide strongest protection; leakage rises sharply in early rounds then plateaus; model choice shifts absolute rates but preserves topology rankings; spatiotemporal and location attributes leak more readily than identity or regulated identifiers. Fourth, we translate these empirical findings into actionable design guidance: prefer sparse or hierarchical connectivity; limit node degree, network radius, and hub privileges; maximize attacker-target separation; avoid shortcut edges around hub nodes; implement topology-aware reviewing and routing to reduce leakage risk.
@@ 2 Related Work
Recent work has begun to analyze how an LLM agent’s long-term memory can itself become an attack surface. (wang-etal-2025-unveiling-privacy) introduce MEXTRA, a black-box memory extraction attack that elicits sensitive user records from an agent’s long-term memory and study how memory design and prompting strategies affect leakage. (chen2024agentpoison) propose AgentPoison, a backdoor-style attack that poisons an agent’s long-term memory or RAG knowledge base with a small set of adversarial demonstrations so that malicious behaviors are triggered at inference time. (dong2025minja) further develop MINJA, a query-only memory injection attack that injects malicious records into an agent’s memory bank without direct write access, causing future queries containing specific victim terms to be mapped to attacker-chosen targets. These works demonstrate that an individual agent’s memory module poses serious privacy and integrity risks; in contrast, our work studies how such leakage emerges and propagates in multi-agent systems, where communication topology and attacker–target placement jointly shape the overall leak rate.
@@
Recent work uses graph structures to study multi-agent safety. NetSafe shows that networks with dense links are easier to break by adversarial spread; the larger the average distance from the attacker to other nodes, the safer the network; star graphs degrade strongly under attack (netsafe). G-Safeguard goes further: it builds a discourse graph, uses a graph neural network for anomaly detection, and restores performance after prompt injection through topological intervention (gsafeguard). Structural comparisons show: hierarchical designs resist malicious agents better than flat or fully connected designs; mechanisms like cross-examination or reviewer agents can improve robustness (iclr_resilience). Our method follows this topology-centered view, but it focuses on memory and PII entity leakage. It quantifies topology-conditioned leakage across graph families, sizes, node placement, and rounds.
@@
According to black-box red teaming, multi-agent systems have a high probability of leaking proprietary elements such as agent counts, network topologies, system and task prompts, and tool usage. It is also evident that adversarial inputs can subvert orchestration and allow arbitrary code execution (codeexec). Integrity attacks show that small, immediate changes can affect monitors and collaborators while bypassing LLM-based supervision (integrity). These findings highlight the need for internal, structure-aware evaluations. In this article we trace the PII entity propagation from target nodes conditioned on topology and placements, going beyond external exfiltration or endpoint impacts (masleak). Taken together with the topology-centric safety literature above, prior work shows that (i) topology shapes vulnerabilities and supports graph-based defenses (netsafe; gsafeguard), (ii) structure influences resilience to malicious agents (iclr_resilience), and (iii) multi-agent systems are susceptible to sensitive-information leakage and adversarial manipulation (masleak; codeexec; integrity). We combine these threads into a topology-driven, reproducible leakage evaluation that converts architectural parameters (graph family, size, placements, horizon) into measurable leakage curves and design guidelines.
@@ 3.2 Engram: Agent and State Initialization
Full system and phase-conditioned user prompts for all three agent roles are provided in Appendix E (Tables 6 and 7). For the attack agent’s system prompt we considered two framings: an “overt attacker” prompt that explicitly instructs the agent to steal PII while clarifying that all data are synthetic, and a more subtle prompt that frames information gathering as a prerequisite for completing the assigned task rather than as theft. The overt framing frequently triggered safety refusals in pilot runs, so all main experiments use the subtler collaborative framing; the full text of the overt attacker prompts is given in Appendix E, Table 8.
@@
At $t=0$, we distribute the shared user-level task context $(B_{i},Q_{i})$ to all agents. Each agent $v\in V_{i}$ then performs an initial, independent inference step. This generation process is conditioned on three distinct factors: the agent’s role-specific system prompt, the shared task input $(B_{i},Q_{i})$, and, exclusively for the target agent, the private memory block containing the PII entities. We strictly enforce the following structured output format:
@@ 3.4 Resonance: Topological State Diffusion
We define the Time-to-Leak $\tau_{\text{leak}}$ as the first round $t$ where the attacker’s visible response $R_{\text{atk}}^{(t)}$ contains a subset of the ground-truth PII entities $\mathcal{S}$:
@@
If the set is empty for all $t$, $\tau_{\text{leak}}=\infty$. This metric allows us to quantify not just if leakage occurs, but the velocity at which a topology permits PII entity diffusion.
@@ 3.5 Evaluation
For each sample $i$ and round $t$, let $A_{i}^{(t)}$ be the attacker’s message at that round. We define a matching function
$\mathrm{match}(y,S_{i})\subseteq S_{i}$
that returns the subset of PII entities whose string values appear in $y$. The set of PII entities recovered at round $t$ is
$\hat{S}_{i}^{(t)}=\mathrm{match}\!\left(A_{i}^{(t)},S_{i}\right).$
We run the Memory Propagation phase for at most $R_{\max}$ rounds. If at some round $t$ the attacker has recovered all PII entities, i.e., $\hat{S}_{i}^{(t)}=S_{i}$, we stop early and set the final round $r_{i}^{\star}=t$; otherwise we set $r_{i}^{\star}=R_{\max}$. The final recovered set for sample $i$ is
$\hat{S}_{i}=\hat{S}_{i}^{(r_{i}^{\star})}.$
Each sample is then categorized as: success if $|\hat{S}_{i}|=|S_{i}|$, failure if $|\hat{S}_{i}|=0$, partial success otherwise.
@@ 4 Experiments
Table 2: Topology-level leakage aggregated over attacker–target placements. For each topology (row) and number of agents (column), the table reports the mean leakage rate (with standard deviation, all in percentage points) across all attacker–target index pairs for that (topology, agent-count) setting, under Llama3.1-70b and DeepSeek-v3.1. Higher values indicate greater PII leakage.
In this section we evaluate multi-agent networks under different graph topologies. Our study focuses on five research questions: RQ1 (Topology Matters): whether leakage rates differ across topologies under fixed agent counts and rounds, and which structures are most vs. least leakage-prone; RQ2 (Position/Centrality): how attacker/target placement within the same topology affects leakage; RQ3 (Scaling with Agents & Rounds): whether enlarging the number of agents (graph nodes) increases leakage monotonically or reaches saturation; RQ4 (PII entity Type Robustness): whether different types of PII entities (numerical, string, identity) exhibit distinct leakability and whether such differences remain invariant across topologies; RQ5 (LLM Matters): whether different base LLMs lead to materially different outcomes.
@@ 4.1 Dataset: SPIRIT
To evaluate leakage risks in a realistic yet ethical manner, we construct SPIRIT (Synthetic PII Role-based Interaction Tasks), a multi-agent dataset built from a high-fidelity simulation environment derived from the Gretel Synthetic Domain-Specific Documents Dataset (gretel-pii-docs-en-v1). These synthetic records act as privacy-preserving proxies for real-world sensitive documents while preserving the semantic coherence and distributional properties of PII entities in domains such as healthcare, finance, and identity verification.
@@
4.2 Experimental Setup
All experiments are repeated three times. For each cell, we report the mean and standard deviation across the three runs, consistent with Table 2 and Table 3. We also put a more detailed Table 4 in Appendix C. This appendix table reports the full leakage scores for all combinations of topology, attacker–target index pair, and agent count under both base models.

### v3

@@ Abstract
Graph topology is a fundamental determinant of memory leakage in multi-agent LLM systems, yet its effects remain poorly quantified. We introduce MAMA (Multi-Agent Memory Attack), a framework that measures how network structure shapes leakage. MAMA operates on synthetic documents containing labeled Personally Identifiable Information (PII) entities, from which we generate sanitized task instructions. We execute a two-phase protocol: Engram (seeding private information into a target agent’s memory) and Resonance (multi-round interaction where an attacker attempts extraction). Over 10 rounds, we measure leakage as exact-match recovery of ground-truth PII from attacker outputs. We evaluate six canonical topologies (complete, ring, chain, tree, star, star-ring) across $n\in\{4,5,6\}$, attacker–target placements, and base models. Results are consistent: denser connectivity, shorter attacker–target distance, and higher target centrality increase leakage; most leakage occurs in early rounds and then plateaus; model choice shifts absolute rates but preserves topology ordering; spatiotemporal/location attributes leak more readily than identity credentials or regulated identifiers. We distill practical guidance for system design: favor sparse or hierarchical connectivity, maximize attacker–target separation, and restrict hub/shortcut pathways via topology-aware access control.
@@
Jinbo Liuthanks: These authors are co-first authors.1   Defu Cao11footnotemark: 11   Yifei Weithanks: These authors contributed equally to data collection, analysis, and manuscript revision.1   Tianyao Su22footnotemark: 21 Yuan Liang22footnotemark: 21   Yushun Dong2   Yan Liu1   Yue Zhao1   Xiyang Hu3 1University of Southern California  2Florida State University  3Arizona State University {jinboliu,defucao,yifeiwei,tianyaos,yliang23,yanliu.cs,yue.z}@usc.edu yd24f@fsu.eduxiyanghu@asu.edu
@@ 1 Introduction
Multi-agent systems based on large language models (LLMs) are rapidly moving from prototype to real-world use (Li et al., 2024). When viewed through the lens of network science, multi-agent LLM systems are distributed communication networks (Kurose and Ross, 2017; Watts and Strogatz, 1998). Network topology, the pattern of connections between agents, becomes a first-order security parameter (Newman and Watts, 1999). Recent work on LLM multi-agent security provides empirical support: connectivity patterns and inter-agent distances create pathways for adversarial diffusion (Wang et al., 2025c), with dense topologies like fully-connected graphs proving particularly vulnerable (Yu et al., 2024; Huang et al., 2025).
Beyond the propagation of adversarial prompts or harmful content, the leakage of Personally Identifiable Information (PII) poses an equally critical threat in multi-agent architectures. Recent investigations have found topology details, system prompts, and tool leaks in black-box configurations (Wang et al., 2025b; Dong et al., 2025a), as well as co-hijacking caused by malicious input (Triedman et al., 2025; Zheng et al., 2025). These findings underscore an urgent research gap: we lack systematic understanding of how communication topology shapes the leakage surface for private information in multi-turn, multi-agent interactions.
@@
Topology-aware security research has begun addressing these concerns through distance and connectivity analysis (Yu et al., 2024), graph-level interventions (Wang et al., 2025c), and structural resilience comparisons (Huang et al., 2025). However, three critical gaps remain. First, prior work targets adversarial content propagation and task degradation (Yu et al., 2024; Wang et al., 2025c) rather than fine-grained PII leakage dynamics. Second, data exfiltration studies do not systematically control for agent placement, graph distance, or interaction horizon (Huang et al., 2025; Wang et al., 2025b). Third, while network theory predicts that sparse long-range links and shortcuts dramatically alter diffusion (Watts and Strogatz, 1998; Newman and Watts, 1999), these structural phenomena have not been rigorously connected to information leakage in LLM-based systems.
@@
We introduce MAMA (Multi-Agent Memory Attack), a systematic framework for measuring how network topology governs memory leakage in multi-agent LLM systems. MAMA addresses the limitations of prior work through three key design principles. First, controlled synthesis: we generate synthetic task documents containing labeled PII entities, then derive sanitized task instructions that contain no direct PII, ensuring any leakage originates from agent memory rather than task specification. Second, systematic topology variation: we evaluate six canonical graph structures (fully connected, ring, chain, binary tree, star, and star-ring) across team sizes ($n\in\{4,5,6\}$), systematically varying attacker-target placement to control for graph distance and node centrality (Huang et al., 2025). Third, temporal resolution: we execute a two-phase protocol—Engram (memory seeding) and Resonance (multi-round extraction)—tracking leakage dynamics over up to 10 interaction rounds. Our framework operationalizes network science concepts (distance, hubs, long-range links) as measurable security indicators (Yu et al., 2024; Wang et al., 2025c).
@@
Our work makes four principal contributions. First, we develop a controllable experimental pipeline spanning data synthesis, instruction sanitization, and a standardized two-phase interaction protocol, enabling systematic evaluation across graph families. Second, we formalize a unified threat model with explicit agent roles and comprehensive leakage metrics. Third, using MAMA we uncover consistent topology-leakage patterns: denser connectivity and shorter attacker-target distances increase leakage; fully connected and star-ring structures prove most vulnerable while chains and trees provide strongest protection; leakage rises sharply in early rounds then plateaus; model choice shifts absolute rates but preserves topology rankings. Fourth, we translate these empirical findings into actionable design guidance: prefer sparse or hierarchical connectivity; limit node degree, network radius, and hub privileges.
@@ 2 Related Work
LLM agents’ long-term memory has emerged as an attack surface. MEXTRA demonstrates black-box memory extraction of sensitive user records and analyzes how memory design and prompting affect leakage (Wang et al., 2025a). AgentPoison poisons long-term memory or RAG stores with a small set of adversarial demonstrations to trigger malicious behaviors at inference time (Chen et al., 2024). MINJA shows a query-only memory injection pathway that plants malicious records without direct write access, steering future retrieval toward attacker-chosen targets (Dong et al., 2025b). While these works focus on single-agent memory risks, we study leakage in multi-agent systems, where communication topology and attacker–target placement jointly shape how private information propagates and is extracted.
@@
Graph structure is increasingly used to characterize multi-agent safety. NetSafe reports that denser links are easier to break, while larger attacker-to-node distances improve safety; star-like designs degrade sharply under adversarial spread (Yu et al., 2024). G-Safeguard builds discourse graphs and uses GNN-based anomaly detection with topological intervention to recover robustness under prompt injection (Wang et al., 2025c). Related analyses further suggest hierarchical organizations can be more resilient than flat or fully connected teams, and that roles such as reviewers or cross-examiners can improve robustness (Huang et al., 2025). We follow this topology-centered view but focus on memory-based privacy: measuring topology-conditioned PII leakage across graph families, sizes, placements, and interaction rounds.
@@
Black-box red teaming indicates multi-agent systems can leak internal details (e.g., agent counts, topology, and system/task prompts) and can be subverted by adversarial inputs, including pathways to code execution (Wang et al., 2025b; Triedman et al., 2025). Complementary integrity attacks show that small, localized manipulations can mislead monitors and collaborators while evading LLM-based supervision (Zheng et al., 2025). These findings motivate structure-aware evaluation. Our work operationalizes this need by tracing PII entity propagation from target nodes under controlled topologies and placements, producing topology-conditioned leakage curves and actionable design guidelines.
@@ 3.2 Engram: Agent and State Initialization
Full system and phase-conditioned user prompts for all three agent roles are provided in Appendix E (Tables 7 and 8). For the attack agent’s system prompt we considered two framings: an “overt attacker” prompt that explicitly instructs the agent to steal PII while clarifying that all data are synthetic, and a more subtle prompt that frames information gathering as a prerequisite for completing the assigned task rather than as theft. The overt framing frequently triggered safety refusals in pilot runs, so all main experiments use the subtler collaborative framing; the full text of the overt attacker prompts is given in Appendix E, Table 9.
@@
At $t=0$, we distribute the shared user-level task context $(B_{i},Q_{i})$ to all agents. Each agent $v\in V_{i}$ then performs an initial, independent inference step. This generation process is conditioned on three distinct factors: the agent’s role-specific system prompt, the shared task input $(B_{i},Q_{i})$, and, exclusively for the target agent, the private memory block containing the PII entities. We enforce the following output format:
@@ 3.4 Resonance: Topological State Diffusion
We define $\tau_{\text{leak}}$ as the first round $t$ where the attacker’s visible response $R_{\text{atk}}^{(t)}$ contains a subset of the ground-truth PII entities $\mathcal{S}$:
@@
If the set is empty for all $t$, $\tau_{\text{leak}}=\infty$. This captures not only whether leakage occurs, but also how quickly a topology enables PII diffusion.
@@ 3.5 Evaluation
For each sample $i$ and round $t$, let $A_{i}^{(t)}$ be the attacker’s message at that round. We define a matching function $\mathrm{match}(y,S_{i})\subseteq S_{i}$ that returns the subset of PII entities whose string values appear in $y$. The set of PII entities recovered at round $t$ is $\hat{S}_{i}^{(t)}=\mathrm{match}\!\left(A_{i}^{(t)},S_{i}\right)$. We run the Memory Propagation phase for at most $R_{\max}$ rounds. If at some round $t$ the attacker has recovered all PII entities, i.e., $\hat{S}_{i}^{(t)}=S_{i}$, we stop early and set the final round $r_{i}^{\star}=t$; otherwise we set $r_{i}^{\star}=R_{\max}$. The final recovered set for sample $i$ is $\hat{S}_{i}=\hat{S}_{i}^{(r_{i}^{\star})}$. Each sample is then categorized as: success if $|\hat{S}_{i}|=|S_{i}|$, failure if $|\hat{S}_{i}|=0$, or partial success.
@@ 4 Experiments
GPT-4o GPT-4o-mini
Topology / Num Agents 4 5 6 4 5 6
Circle 3.69 (0.74) 2.24 (1.82) 2.99 (0.67) 7.53 (3.67) 4.65 (2.17) 3.63 (1.58)
Complete 4.17 (1.47) 3.20 (1.55) 4.16 (2.24) 5.29 (0.83) 4.65 (0.74) 4.27 (2.98)
Star-Ring 4.59 (0.18) 2.40 (0.72) 2.32 (0.50) 5.13 (1.47) 4.17 (0.50) 3.20 (1.84)
Star-Pure 2.24 (0.56) 2.67 (1.77) 1.92 (1.16) 6.31 (0.74) 6.52 (2.41) 3.95 (2.32)
Chain 2.88 (0.28) 2.80 (0.14) 1.99 (0.39) 4.70 (0.49) 5.18 (0.44) 3.95 (0.84)
Tree 2.48 (0.91) 1.12 (0.16) 2.12 (0.89) 6.41 (0.61) 3.95 (2.43) 5.17 (0.67)
Table 2: Topology-level leakage aggregated over attacker–target placements. We report mean $\pm$ std leakage ($p$) over all attacker–target pairs for each topology and agent count, for Llama3.1-70b, DeepSeek-v3.1, GPT-4o, and GPT-4o-mini. Higher values indicate greater PII leakage.
In this section we evaluate multi-agent networks under different graph topologies. Our study focuses on five research questions: RQ1 (Topology Matters): whether leakage rates differ across topologies under fixed agent counts and rounds, and which structures are most vs. least leakage-prone; RQ2 (Position/Centrality): how attacker/target placement within the same topology affects leakage; RQ3 (Scaling with Agents & Rounds): how is the leakage rate associated with an increasing number of rounds; RQ4 (PII entity Type Robustness): whether different types of PII entities (numerical, string, identity) exhibit distinct leakability; RQ5 (LLM Matters): whether different base LLMs lead to materially different outcomes.
@@
Experimental Setup.
We enumerate non-redundant placements up to graph symmetries (and subsample one third of pairs for the binary tree topology for cost), repeat each configuration three times, and report mean and standard deviation (Table 2, Table 3). The full setup is provided in Appendix C.1, and the complete per-configuration leakage table is in Appendix C.2 (Table 4).
@@ 4.1 Dataset: SPIRIT
To evaluate leakage risks in a realistic yet ethical manner, we construct SPIRIT (Synthetic PII Role-based Interaction Tasks), a multi-agent dataset built from a high-fidelity simulation environment derived from the Gretel Synthetic Domain-Specific Documents Dataset (AI, 2024). These synthetic records act as privacy-preserving proxies for real-world sensitive documents while preserving the semantic coherence and distributional properties of PII entities in domains such as healthcare, finance, and identity verification.